A boy of mass 60kg is standing over a platform of mass40 kg placed over smooth horizontal surfaceHe throws a stone of mass of 1kg with vel 10ms at an angle of 45 with respect to groundFind the displacement of platform with boy on the horizontal surface when the stone lands on the ground (take g=10mss)

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

Let v be velocity of boy+platform. from conservation of horizontal linear momentum, 100v=10/√2, v=√2/10=0.1414m/s.

time of flight of stone=t, 10/√2=gt/2,t=√2 sec,displacement of platform =vt=(√2/10)*√2= 0.2 m